# MemAscend: System Memory Optimization for SSD-Offloaded LLM Fine-Tuning

### v3

@@ II-A SSD offloading framework
In ZeRO-Infinity’s workflow, forward and backward passes are executed on the GPU, while the optimizer step is typically handled on the CPU due to its lower computational complexity. This is because transferring the optimizer step to the GPU is often not worthwhile, given the overhead involved [32]. To optimize data movement, ZeRO-Infinity employs techniques such as bandwidth-centric partitioning and overlap-centric design, which aim to improve transfer efficiency across the PCIe and NVMe interfaces. While bandwidth-centric partitioning partitions one large tensor and stagger multiple tensors to fully saturate available NVMe bandwidth during each transfer window, the overlap-centric design asynchronously loads the next set of weights or gradients from SSDs into system memory during forward passes executing on GPUs. Despite its effectiveness in scaling model size, ZeRO-Infinity’s reliance on CPU system memory as a staging buffer between the GPU and SSD can lead to significant inefficiencies. These overheads become particularly pronounced on resource-constrained systems with limited system memory, an issue that this work seeks to address.
@@ III-A System Memory Fragmentation Issue
In the SOTA SSD offloading technique, a system memory buffer pool, referred to as the parameter buffer pool, is designed to prefetch model weights from SSDs into DRAM for a brief window while each layer executes its forward or backward pass. Nonetheless, the current design of the system memory buffer pool leads to the issue of system memory fragmentation that inflates memory consumption well beyond what the models truly need. The current design equips the buffer pool with pre-allocated pinned memory buffers of the same size. Then, when a layer requires its weights, the buffer pool assigns a pinned memory buffer to it. The weights are loaded from SSD into this memory buffer via I/O operations, followed by a host-to-device copy to transfer the data to the GPU. Once the weights are no longer needed, the memory buffer is returned to the pool. Since prefetching is employed, multiple transformer blocks [^2] Transformer blocks are the fundamental building units of the transformer architecture and are designed to process input sequences by capturing relationships between elements through attention mechanisms and feed-forward networks. are typically in flight at once. The required buffer pool size, therefore, equals the size of each memory buffer multiplied by the number of memory buffers needed per transformer block, and further multiplied by the number of transformer blocks in flight.
@@ III-D Filesystem Overhead in SSD offloading
In an SSD offloading system, the SSD serves as extended memory. It stores low-precision compute weights, such as fp16, along with the full-precision fp32 master weights, gradients, and optimizer states. These data are loaded or offloaded depending on the computational context. The current implementation uses conventional file systems with the O-DIRECT flag, which bypasses the page cache to ensure that data transfers go directly to the SSD. Each tensor is offloaded to a separate file, allowing file systems, such as ext4, to manage storage allocation without direct intervention. However, writing to ext4 carries overhead. The file system maintains metadata for each file, including name, size, permissions, and timestamps, which requires additional I/O and increases latency. Allocating disk blocks can also involve searching for contiguous space, updating allocation tables, and maintaining alignment, all of which add computational cost and risk fragmentation. When the SSD is used as an extension of memory and subjected to sustained I/O, these file system overheads accumulate. This added latency can limit the efficiency gains expected from offloading, particularly during training.
@@ IV-B Adaptive Buffer Pool
Notably, the management cost of the adaptive buffer pool is kept minimal by following the same underlying allocation approach as ZeRO-Infinity. This approach first allocates a single large, sequential monolithic tensor and then maintains metadata to manage which portion is assigned to each buffer. This design minimizes the allocation overhead that might occur with multi-pool management. When a buffer request occurs, such as when a tensor needs to be fetched from the SSD and copied to the CPU, a hashtable records a unique identification key. This key maps to the data structure containing the buffer’s metadata, which includes its size and the index of its portion within the large monolithic tensor. Since the number of distinct buffers is typically small (fewer than 100 in most mainstream open-source models, depending on the configuration), the complexity overhead of managing this hashtable metadata is negligible in terms of both computation and memory usage. Furthermore, this scheme efficiently generalizes to multi-GPU deployments that use parameter partitioning. In such a setup, each process loads only its own parameter shard, causing the per-process buffers to shrink proportionally with the number of partitions. Thus, the adaptive buffer pool maintains its efficiency in both single-GPU and distributed configurations.
@@ IV-C Alignment Free Pinned Memory Allocation
As noted in Section III-B, PyTorch’s standard caching allocator applies a power-of-two alignment policy. This policy introduces substantial memory waste when managing the large and static tensors essential for SSD offloading. These tensors are allocated once at initialization and persist until training completes, and they are major contributors to overhead when the power-of-two alignment is enforced due to their large size. These critical, parameter-related buffers include: (1) the gradient flat buffer for holding GPU-computed gradients during the backward pass, (2) all buffers within the parameter buffer pool used for prefetching model parameters from SSD to GPU, and (3) the optimizer state buffer that facilitates fetching and updating optimizer state subgroups on the CPU. To mitigate memory waste in these buffers, MemAscend incorporates a custom C++ extension to bypass this behavior and ensure memory efficiency through a precise, three-step process.
Direct Memory Allocation. First, the extension directly allocates CPU memory using posix_memalign. Unlike PyTorch’s default allocator, it aligns memory only to a 4096-byte boundary, which is the minimum required for Direct Memory Access (DMA) operations, thus eliminating the overhead from power-of-two rounding. Page-Locking and CUDA Registration. Next, the allocated memory is page-locked and registered with all available CUDA contexts via the cudaHostRegister API, using the cudaHostRegisterPortable flag. This step makes the memory block accessible to all GPUs for high-speed transfers. Pinned Tensor Creation. Finally, a CPU-pinned tensor is created from this memory block using the torch::from_blob API. This allows the PyTorch engine to use the buffer for efficient GPU DMA transfers without additional memory copies or the need to involve the CPU.
@@ VI-B2 Assessment of the Integrated Solution
To evaluate the overall memory efficiency of the proposed MemAscend, a comparison of peak system memory usage against the baseline ZeRO-Infinity system across four model architectures is conducted. This comparison is illustrated in Figure 15, a bar graph designed to highlight the system memory reduction achieved by our approach.
The evaluation results reveal a consistent trend. MemAscend significantly reduces system memory usage across all models tested compared to the baseline. Specifically, for the Llama3.1-8B model, the baseline consumes 91.06 GiB, while our method reduces this to 44.71 GiB, a savings of approximately 50.9%. For the Qwen2.5-7B model, memory usage drops from 109.06 GiB to 43.67 GiB, leading to a reduction of 60.0%. In the Qwen2.5-14B case, the baseline requires 174.5 GiB, whereas MemAscend uses only 76.1 GiB, achieving a 56.4% decrease. Finally, for the Qwen2.5-32B model, the baseline demands 322.3 GiB, while our method lowers this to 143.6 GiB, resulting in a reduction of 55.4%. On average, MemAscend achieves a memory reduction of approximately 55.7% across these models. This substantial memory savings underscores the effectiveness of our approach in optimizing resource utilization, with free memory allowing for further enhancements, such as supporting a longer context by utilizing the released space for activation checkpoint values, as discussed in the following sections. The effectiveness of MemAscend is the result of a combination of optimizations detailed in Section V-A.

### v4

@@ II-A SSD offloading framework
ZeRO-Infinity executes forward and backward passes on the GPU, but it typically offloads the optimizer step to the CPU. This is because the optimizer’s low arithmetic intensity rarely justifies the PCIe transfer costs required to move optimizer states back to the GPU [32]. DeepSpeed supports this via a host optimizer with a fused C++/AVX backend. To maximize efficiency, this backend uses contiguous tensors and template dispatch for kernel selection. It supports fp16/bf16 parameters paired with either fp32 or low-precision optimizer states. For performance, updates are vectorized using AVX512/AVX2, while an OpenMP-parallel tiled loop handles remaining elements, including bias correction and decoupled weight decay. To optimize PCIe/NVMe throughput, ZeRO-Infinity employs bandwidth-centric partitioning and overlap-centric execution. By chunking tensors and staggering transfers, the system saturates NVMe bandwidth while asynchronously staging upcoming weights from the SSD to DRAM. However, relying on CPU memory as a staging buffer remains a bottleneck on systems with limited DRAM, a limitation this work aims to resolve.
@@ III-A System Memory Fragmentation Issue
State-of-the-art (SOTA) SSD offloading techniques utilize a parameter buffer pool in system memory to prefetch weights from SSDs during layer execution. However, current designs suffer from memory fragmentation, inflating consumption beyond actual model needs. This occurs because the pool relies on pre-allocated, uniform-sized pinned buffers. When a layer requests weights, it occupies a buffer for SSD-to-Host I/O and subsequent Host-to-Device (H2D) transfer before returning it to the pool. Since prefetching keeps multiple transformer blocks [^2] Transformer blocks are the fundamental building units of the transformer architecture and are designed to process input sequences by capturing relationships between elements through attention mechanisms and feed-forward networks. "in flight" simultaneously, the total pool size scales as the product of the buffer size, the buffers required per block, and the number of blocks in flight. Even though GPUDirect Storage (GDS) bypasses system DRAM, the core issue of inefficient buffer management remains. While ZeRO-Infinity uses GDS to load weights directly from NVMe SSDs into GPU memory, it merely relocates the buffer pool from system memory to GPU memory. In resource-constrained fine-tuning, this reserved GPU capacity competes with activations and intermediates, directly limiting batch size, context length, and model scale. Consequently, this work focuses on the system memory setting to address these allocation inefficiencies.
@@ III-D Filesystem Overhead in SSD offloading
In an SSD offloading system, the SSD serves as extended memory. It stores low-precision compute weights, such as fp16, along with the full-precision fp32 master weights, gradients, and optimizer states. These data are loaded or offloaded depending on the computational context. The current design, also known as DeepNVMe in ZeRO Infinity, uses conventional file systems with the O_DIRECT flag, which bypasses the page cache to ensure that data transfers go directly to the SSD. Each tensor is offloaded to a separate file, allowing file systems, such as ext4, to manage storage allocation without direct intervention. However, writing to ext4 carries overhead. The file system maintains metadata for each file, including name, size, permissions, and timestamps, which requires additional I/O and increases latency. Allocating disk blocks can also involve searching for contiguous space, updating allocation tables, and maintaining alignment, all of which add computational cost and risk fragmentation. When the SSD is used as an extension of memory and subjected to sustained I/O, these file system overheads accumulate. This added latency can limit the efficiency gains expected from offloading, particularly during training.
@@ IV-B Adaptive Buffer Pool
Notably, the management cost of the adaptive buffer pool is kept minimal by following the same underlying allocation approach as ZeRO-Infinity. This approach first allocates a single large, sequential monolithic tensor and then maintains metadata to manage which portion is assigned to each buffer. This design minimizes the allocation overhead that might occur with multi-pool management. When a buffer request occurs, such as when a tensor needs to be fetched from the SSD and copied to the CPU, a hashtable records a unique identification key. This key maps to the data structure containing the buffer’s metadata, which includes its size and the index of its portion within the large monolithic tensor. Since the number of distinct buffers is typically small (fewer than 100 in most mainstream open-source models, depending on the configuration), the complexity overhead of managing this hashtable metadata is negligible in terms of both computation and memory usage. Furthermore, this scheme efficiently generalizes to multi-GPU deployments that use parameter partitioning. In such a setup, each process loads only its own parameter shard, causing the per-process buffers to shrink proportionally with the number of partitions. Notably, the adaptive buffer pool extends naturally to GDS-based offloading. Since GDS necessitates similar buffer management on the GPU, this method provides a robust framework for minimizing VRAM overhead, demonstrating its broad applicability across different memory tiers.
@@ IV-C Alignment Free Pinned Memory Allocation
As noted in Section III-B, PyTorch’s standard caching allocator applies a power-of-two alignment policy. This policy incurs substantial memory waste when managing the large, static tensors essential for SSD offloading. These tensors are allocated once at initialization and persist until training completes, and they are a major contributor to overhead when power-of-two alignment is enforced due to their large size. These critical, parameter-related buffers include: (1) the gradient flat buffer for storing GPU-computed gradients during the backward pass, (2) the parameter buffer pool for prefetching SSD-resident parameters, and (3) the optimizer state buffer for fetching and updating optimizer state subgroups on the CPU. To mitigate this waste, MemAscend uses a custom C++ extension to bypass PyTorch’s default alignment.
Direct Memory Allocation: The extension uses posix_memalign to allocate CPU memory with 4096-byte alignment, satisfying DMA requirements while eliminating power-of-two rounding overhead. Page-Locking and CUDA Registration: Allocated regions are page-locked and registered via cudaHostRegister with the cudaHostRegisterPortable flag, enabling high-throughput DMA transfers across all CUDA contexts. Pinned Tensor Creation and Lifecycle Management: MemAscend wraps the memory into a torch::from_blob tensor with a custom deleter that leverages PyTorch’s reference counting to safely trigger cudaHostUnregister and free exactly once, bypassing the CPU caching allocator to prevent bookkeeping interference and memory leaks.
@@ VI-B2 Assessment of the Integrated Solution
To evaluate the overall memory efficiency of the proposed MemAscend, a comparison of peak system memory usage against the baseline ZeRO-Infinity system across four model architectures is conducted. This comparison is illustrated in Figure 15, a bar graph designed to highlight the system memory reduction achieved by our approach. The evaluation results reveal a consistent trend. MemAscend significantly reduces system memory usage across all models tested compared to the baseline. Specifically, for the Llama3.1-8B model, the baseline consumes 91.06 GiB, while our method reduces this to 44.71 GiB, a savings of approximately 50.9%. For the Qwen2.5-7B model, memory usage drops from 109.06 GiB to 43.67 GiB, leading to a reduction of 60.0%. In the Qwen2.5-14B case, the baseline requires 174.5 GiB, whereas MemAscend uses only 76.1 GiB, achieving a 56.4% decrease. Finally, for the Qwen2.5-32B model, the baseline demands 322.3 GiB, while our method lowers this to 143.6 GiB, resulting in a reduction of 55.4%. On average, MemAscend achieves a memory reduction of approximately 55.7% across these models. This substantial memory savings underscores the effectiveness of our approach in optimizing resource utilization, with free memory allowing for further enhancements, such as supporting a longer context by utilizing the released space for activation checkpoint values, as discussed in the following sections. The effectiveness of MemAscend is the result of a combination of optimizations detailed in Section V-A.